OpenStudy (anonymous):

A polynomial f(x) with real coefficients and leading coefficient 1 has zeros -7, 0, 1 + i and degree 4. Express f(x) as a product of linear and quadratic polynomials with real coefficients that are irreducible over IR. (Another one, I tried but it didn't work)

OpenStudy (anonymous):

f(x)= x(x+7)(x -1- i)(x -1 +i ) f(x) = x(x+7)[(x-1)^2 - i^2] f(x)= x(x+7)[(x-1)^2 +1) f(x)= x(x+7)(x^2 -2x +1 +1) f(x) x(x+7)(x^2 -2x +2)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

its probably a lack of brackets in your working out that is causing the mistakes lol

OpenStudy (anonymous):

Probably :)

OpenStudy (anonymous):

yes, bracket is pretty important here :D

OpenStudy (anonymous):

if 1+i is a root then so is 1-i

OpenStudy (anonymous):

f(x) = z (z-+7)(z - (1+i) ) (x- (1-i) )

OpenStudy (anonymous):

you probably need to put the brackets inside brackets , im guessing you are forgetting a minus sign when expanding

OpenStudy (anonymous):

why did I put z and x :|

OpenStudy (anonymous):

Yep, that's where it went wrong!

OpenStudy (anonymous):

lol lol, z and x...variables are always painful, we don't want more of that :D